OpenStudy (anonymous):

On Sunday, a local hamburger shop sold a combined total of 375 hamburgers and cheeseburgers. The number of cheeseburgers sold was two times the number of hamburgers sold. How many hamburgers were sold on Sunday?

OpenStudy (shamil98):

C = 2h h = c/2 C + H = 375 2h + h = 375 3h = 375 h= 125